# Detecting Document-level Paraphrased Machine Generated Content: Mimicking Human Writing Style and Involving Discourse Features

### v1

@@ Abstract
The availability of high-quality APIs for Large Language Models (LLMs) has facilitated the widespread creation of Machine-Generated Content (MGC), posing challenges such as academic plagiarism and the spread of misinformation. Existing MGC detectors often focus solely on surface-level information, overlooking implicit and structural features. This makes them susceptible to deception by surface-level sentence patterns, particularly for longer texts and in texts that have been subsequently paraphrased.
To overcome these challenges, we introduce novel methodologies and datasets. Besides the publicly available dataset Plagbench, we developed the paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) datasets using GPT and DIPPER, a discourse paraphrasing tool, by extending artifacts from their original versions. To address the challenge of detecting highly similar paraphrased texts, we propose MhBART, an encoder-decoder model designed to emulate human writing style while incorporating a novel difference score mechanism. This model outperforms strong classifier baselines and identifies deceptive sentence patterns. To better capture the structure of longer texts at document level, we propose DTransformer, a model that integrates discourse analysis through PDTB preprocessing to encode structural features. It results in substantial performance gains across both datasets – 15.5% absolute improvement on paraLFQA, 4% absolute improvement on paraWP, and 1.5% absolute improvemene on M4 compared to SOTA approaches.
Detecting Document-level Paraphrased Machine Generated Content: Mimicking Human Writing Style and Involving Discourse Features
Yupei Li1, Manuel Milling2, Lucia Specia1, Björn W. Schuller1,2, 1GLAM team, Imperial College London, London, UK, 2CHI – Chair of Health Informatics, Technical University of Munich, Munich, Germany,
1 Introduction
As LLMs such as GPT-4 OpenAI (2023) and Llama3 Touvron et al. (2023) develop rapidly and become widely utilized by the general public, Machine-Generated Content (MGC) has become increasingly abundant. In particular, people often rely on these models to revise long texts, such as essays, which results in paraphrased MGC. One example is shown in Figure 1, where the original human-written text is highlighted in red on the left side, and the MGC in green on the right side, with differences shaded. It can be argued that content features such as fluency, readability, and clarity all improve after revision, and that the usage of LLM-usage can be hard to detect manually.
@@ 1 Introduction
There are well known risks associated with the misuse of MGC. Crothers et al. (2023) conducted a survey on the applications of MGC, reporting that it may contribute to academic misconduct, the spread of fake political news, and an increase in social media spam. Our goal in this paper is the effective detection of document-level paraphrased MGC, which is rewritten by LLMs from an original human written text.
Automatically distinguishing MGC from Human-Produced Content (HPC) is, however, very challenging, given that MGC tends to resemble HPC, often containing at least chunks that are reproduced from human produced training data. Currently, methods to detect MGC involve a combination of processes, namely a Natural Language Understanding (NLU) module that extracts linguistic and contextual features, followed by a classification module which fits this to binary labels using algorithms such as Logistic Regression (LR) Cox (1958) for decision-making.
However, these suffer from various limitations. First, current NLU models built using Transformers Vaswani et al. (2017) such as Roberta Liu et al. (2019) demonstrate good performance in analysing short texts (such as sentences) but less so when processing document-level data, as shown in Fu (2024). They tend to struggle with retaining information if the texts are long. Even LLMs are limited in the length of the input texts that they can process effectively. Document-level texts exhibit more complex structures’ relationships. Secondly, the degree of similarity between MGC and HPC is higher especially when LLMs are used to paraphrase HPC rather than generating it from scratch. This makes distinguishing HPC and MGC challenging, as differences in linguistic content are less distinctive. Previous work focuses on identifying similarities at surface level linguistic features, ignoring the importance of other implicit features such as writing style or structural features, which we show can contribute to the effectiveness of a detection model. There is indeed a difference in writing style in Figure 1: the left side is more poetic, while the right side is a plainer narrative, though the distinction may not be obvious to the human reader. These limitations lead to underperforming detectors in work to date.
We propose two novel approaches to automatically detect MGC which address the limitations described above. The aim is to better classify the texts by identifying structural features, and implicit features, such as discourse relations or writing styles. Specifically, we introduce two models: MhBART, which does detection through writing style by leveraging the capabilities of BART Lewis et al. (2019), and DTransformer, which incorporates discourse features to improve detection accuracy.
@@
A. We create paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) – two parallel HPC-MGC corpora for model evaluation and further research in Section 3
B. We introduce MhBART – an encoder-decoder model designed to learn human writing style, which outperforms strong baselines for paraphrased MGC detection in Section 4. Additionally, we identified sentence patterns that can deceive current detectors.
C. We introduce DTransformer – a discourse-aware Transformer-based model to incorporate discourse features in detection, which leads to better performance for long form texts in Section 5.
2 Related work
2.1 MGC Detection Approaches
@@
In the academic literature, two types of models are commonly used: statistical methods, also known as feature-based methods, and neural network methods. Herbold et al. (2023) identified seven linguistic features and criteria for distinguishing MGC from HPC, and they concluded that MGC tends to exhibit fewer discourse and epistemic words but displays greater lexical diversity and nominalizations. Choudhary and Arora (2021) and GLTR (Generative Language Test Room) Gehrmann et al. (2019) also made decision with linguistic features.
@@
However, feature-based methods need manual work to design the features. There have been studies comparing neural network methods and statistical methods Hayawi et al. (2023), demonstrating the superior efficiency of Transformer-based models Vaswani et al. (2017), such as Roberta Large OpenAI Detector Solaiman et al. (2019), the first fake news detector Grover Zellers et al. (2019), a zero-shot detection method DetectGpt Mitchell et al. (2023), watermarking based detection Fu et al. (2024). The Roberta Large Detector has been fined-tuned with 1.5B output of GPT2. The detection is treated as a binary classification. However, it was trained on texts produced with GPT-2, which may retain biases.
@@
Additionally, Fu et al. (2024) have introduced semantic-aware watermarking algorithms, while Jiang et al. (2023) employ perturbation strategies to evade detection methods based on watermarking. Besides, Hu et al. (2023) introduced another adversarial learning approach named Radar, noted for its robust transferability. Additionally, Chaka (2023) directly utilizes LLMs Brown et al. (2020) for MGC detection, but concludes that these models are currently inadequate. Conversely, Chen et al. (2023) combine two LLMs, BERT Devlin et al. (2019), and T5 Raffel et al. (2020), to improve detection efficacy. Moreover, recent advancements include the M4 model for multi-modality MGC detection Wang et al. (2023) and DeBERTa He et al. (2020) for monitoring mainstream news websites Hanley and Durumeric (2024).
Moreover, numerous widely utilized commercial tools commercial tools include GPTZero  [^1] https://gptzero.me/, Copyleaks [^2] https://copyleaks.com/ai-content-detector, Content at Scale  [^3] https://contentatscale.ai/ai-content-detector/, Originality  [^4] https://originality.ai/, Winston AI  [^5] https://gowinston.ai/, and Hive  [^6] https://hivemoderation.com/ai-generated-content-detection. However, these tools utilize models that are not publicly accessible, which makes understanding these tools challenging.
As discussed above, the open challenges that have not covered in previous work are as follows: Combining implicit features with linguistic features, and detecting document-level paraphrased machine-generated texts.
2.2 Detection of Paraphrased-MGC
MGCs learn from human texts which have similar features as HPCs, so MGCs and HPCs become more similar, especially for machine-paraphrased MGC datasets. Krishna et al. Krishna et al. (2023) developed an attack method targeting recent detectors by employing paraphrased sentences in MGC, resulting in a significant decrease in detector accuracy. This approach utilized paraphrased sentence patterns introduced by an external model known as DIPPER Krishna et al. (2023), highlighting the challenges associated with detecting discourse-paraphrased sentences.
@@ 2.2 Detection of Paraphrased-MGC
Efforts have been made to address the paraphrased MGC detection issue. Wahle et al. (2022) and Foltỳnek et al. (2020) proposed using pre-trained word embedding models for text representation, combined with classifiers, to predict paraphrased plagiarism in writing. Animesh et al.  Nighojkar and Licato (2021) introduced an adversarial method to detect paraphrasing by leveraging the inferential properties of sentencesOther research has focused on developing corpus-based methods for detecting paraphrasing Vrbanec and Meštrović (2020). However, paraphrasing involves more than just substituting words and should incorporate additional detectable features. Researchers have manually crafted features, as discussed by Shahmohammadi et al.  Shahmohammadi et al. (2021), which align with syntactic similarity, and other factors.
The aforementioned methods primarily focus on content similarity detection. However, these sentences having formats of discourse paraphrasing are hard to be addressed, which involve changing the structure of sentences and the organization of paragraphs.
2.3 Detection of Document-level MGC
@@
Long texts have more complex language organization compared to shorter texts, so comprehending long texts is more challenging than shorter ones. Current Transformer-based models can fail when processing longer sequences Wolf et al. (2020) and there remains a risk of information loss due to the constrained memory and computational costs associated with self-attention operations Xie et al. (2023).
@@
To address the aforementioned issue, hierarchical models have been proposed that handle sentence-level information before passing on these representations for document-level processing. Zhang et al. (2019) proposed the HIBERT model for document summarization. Moreover, Xie et al. (2023) divided long sequences into different blocks, while SPECTER Cohan et al. (2020) was proposed to embed research papers by utilizing only the metadata. Longformer Beltagy et al. (2020) introduced a sparse attention mechanism that adapts to sequence length, and Grail et al. (2021) proposed to globalize sequences processed by BERT using a bidirectional GRU network to propagate information. Each of these models aim to address memory limitations and computational complexity by selecting representative embeddings. However, there are no dominating models utilizing discourse features.
@@
The discourse features could be represented by two annotations: Rhetorical Structure Theory Discourse Treebank (RST-DT) Carlson et al. (2006), a dataset that captures discourse relations within text spans, represented in a tree structure and the Penn Discourse Treebank (PDTB) Prasad et al. (2008), offering a straightforward representation of relations, specifying two arguments along with their relations Webber et al. (2019) (as depicted in Figure 2). Annotated by human experts, PDTB’s 3.0 version includes over 100 types of relations, demonstrating exceptional applications when combined with Bert models Kim et al. (2020). These annotated labels are detailed about the role of each sentences as well.
@@
Figure 2: PDTB annotation example Lin et al. (2010): It shows the implicit relations, ACCORDINGLY, between the given two sentence.
3 Datasets
@@
Currently, limited public datasets are available that focus on document-level paraphrased texts for MGC detection. For example, the recently published adversarial perturbations dataset introduced revised MGC texts with word replacements Huang et al. (2024), but the dataset does not include meaningful differences in writing style between the two versions. We present the Plagbench Lee et al. (2024) and Multi-domain, and Multi-lingual Black-Box Machine-Generated Text Detection (M4) dataset Wang et al. (2024) as publicly available datasets, and also introduce two new dataset that we created based on a portion of well-known datasets used in other NLP tasks: LFQA Fan et al. (2019), which was also used in the DIPPER dataset Krishna et al. (2023) and WritingPrompts (WP) Fan et al. (2018).
@@
The Plagbench dataset Lee et al. (2024) is generated by GPT specifically for the task of plagiarism detection. We select the paraphrasing plagiarism subset, which consists of 135 data pairs.
@@
The M4 dataset contains a diverse array of human-written sources, including document-level texts from WikiHow, and others. It also includes MGC versions produced by a wide range of generators, such as BLOOMz Team (2022) and ChatGPT. These machine-generated texts often differ in meaning from the original sources because the prompts used are not direct paraphrases but involve expansions, such as the prompt “Write a peer review based on the article." It has been extensively utilized as a benchmark in MGC detection research. The dataset has a total of 133,551 text samples.
@@
For the paraLFQA dataset, the original answers from LFQA are considered as HPC. MGC was produced using the DIPPER model from the HPCs, resulting in discourse-paraphrased-style sentences, which we present as a new dataset. In accordance with the methodology outlined in the original paper, we present T5 as the tokenizer. We selected a lexicon code of 40 and a re-order code of 100. These values represent that 40% of the lexicon and 100% of the order have been modified. These hyperparameters were chosen to emphasize changes in the organization of the language within the dataset. The dataset has around 2,000 MGC-HPC paragraph pairs with around 250-350 words in each paragraph. This serves as a control dataset where the semantic features are similar between the paraphrased and original content, but the writing style is different.
@@
For the paraWP dataset, human-written stories from the provided prompts are considered as HPC. The MGC paraphrases are produced by GPT-3.5-turbo with the following prompt: “Summarize the text. Expand the idea into new text. Keep the original narrative, such as he, she, I, or their names. Return the same number of words as the input texts." In this way, the data has one-to-one matched text pairs with HPC and MGC, which are regarded as paraphrased-style versions. We have created 166,247 documents both for HPC and MGC in the training set, and 12,800 and 12,316 documents for the validation and test sets, respectively. We used BPE tokenizer to process this data and the vocabulary size of the BPE modelling is 8,016. We have also investigated linguistic features of the dataset in Appendix A.
@@
4 MhBART: Mimicking human writing style Bidirectional and Auto-Regressive Transformers
This section introduces our first method especially for the first challenge: detecting MGC with high similarity to the original text. MhBART focuses on learning to mimic human writing style, aiming to detect MGC by identifying deviations from natural language patterns typically found in human-authored content. The method reverses the process of generating MGC in LLMs like GPT, i. e., our model is learning to generate HPC from MGC, such that we can implicitly grasp the writing style of humans.
The motivation behind this approach lies in the capability of language models to comprehend language, enabling them to learn how humans organize words and their overall written style, which constitutes the implicit features of the text. By training the model to transfer from MGC to HPC, the expectation is that it could capture and replicate human-like writing patterns and style effectively. We selected BART Lewis et al. (2019) as the base pre-trained model due to its ability to handle long texts.
The overall model architecture is illustrated in Figure 3. It learns from pairs of MGC and HPC samples that have nearly identical content and ideas but differ in their organization or style. For training, the input to the MhBART consists of a document from this MGC-HPC paired set, and the target (or supervision) is the corresponding HPC. This setup is reminiscent of a translation model, where the source-side sentences are MGC and the target-side sentences are HPC.
During inference and classification, the trained model evaluates the difference between inputs and outputs using various string similarity metrics. Specifically, we employ BLEU Papineni et al. (2002), which assesses similarities in form at the n-gram level, and cosine similarity score, which measures similarity of representations in a shared embedding space and assesses more features based on what the embedding represents. We calculate the cosine similarity score based on the representations of sentences obtained from BERT. We also investigate the cosine similarity score calculated based on universial authorship representation (UAR) Soto et al. (2021) but it shows quite negative results (details discussed in Appendix A). We deem this phenomenon is attributed to the limitations of UAR in effectively capturing the entire writing style and its challenges in fine-tuning for our specific task, so we did not utilize that. We use the four components of the BLEU score and the cosine similarity score as features for a logistic regression (LR) classification model. This LR model learns the boundary conditions necessary to classify texts as either HPC or MGC.
The hypothesis is that BART, when trained to generate texts that adhere to human-written style, will produce outputs that differ significantly from input texts if they are MGC as the model is trained to learn HPC writing style. Therefore, if the model receives HPC input, it would not change the writing significantly.
Figure 3: MhBART approach: Learning to translate from MGC to HPC. The model emulates HPC writing style by training with BART with the input being the corresponding MGC. An intermediate output, which captures properties indicative of human writing style, is generated by the trained BART model and subsequently compared with the input text. A substantial difference score suggests the text is MGC, whereas a low difference score supports classification as HPC.
5 DTransformer
This section introduces our second method, specifically designed to address the second challenge: detecting document-level MGC texts paraphrased from HPC. The DTransformer model focuses on integrating structural features with semantic features, aiming to detect MGC by identifying the ways humans and machines organise text blocks, such as paragraphs. This model is inspired by the discourse information extraction method with PDTB annotation and the effects of structural features discussed in the Multi-Method Qualitative Text Framework Alejandro and Zhao (2023). The model is depicted in Figure 4, with detailed sub-modules discussed in the following paragraphs.
Figure 4: DTransformer model: it incorporates both structural and semantic features. It first splits the paragraphs and employs a hierarchical model to capture document semantic features. Then, using a cross-attention mechanism, it analyzes discourse features for improved classification.

### v2

# Discourse Features Enhance Detection of Document-Level Machine-Generated Content

@@ Abstract
The availability of high-quality APIs for Large Language Models (LLMs) has facilitated the widespread creation of Machine-Generated Content (MGC), posing challenges such as academic plagiarism and the spread of misinformation. Existing MGC detectors often focus solely on surface-level information, overlooking implicit and structural features. This makes them susceptible to deception by surface-level sentence patterns, particularly for longer texts and in texts that have been subsequently paraphrased. To overcome these challenges, we introduce novel methodologies and datasets. Besides the publicly available dataset Plagbench, we developed the paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) datasets using GPT and DIPPER, a discourse paraphrasing tool, by extending artifacts from their original versions. To better capture the structure of longer texts at document level, we propose DTransformer, a model that integrates discourse analysis through PDTB preprocessing to encode structural features. It results in substantial performance gains across both datasets – 15.5% absolute improvement on paraLFQA, 4% absolute improvement on paraWP, and 1.5% absolute improvemene on M4 compared to SOTA approaches. The data and code are available at this link [^1] https://github.com/myxp-lyp/Discourse-Features-Enhance-Detection-of
-Document-Level-Machine-Generated-Content.git
I Introduction
As LLMs such as GPT-4 [1] and Llama3 [2] develop rapidly and become widely utilized by the general public, Machine-Generated Content (MGC) has become increasingly abundant. In particular, people often rely on these models to revise long texts, such as essays, which results in paraphrased MGC. One example is shown in Figure 1, where the original human-written text is highlighted in red on the left side, and the MGC in green on the right side, with differences shaded. It can be argued that content features such as fluency, readability, and clarity all improve after revision, and that the usage of LLM-usage can be hard to detect manually.
@@ I Introduction
There are well known risks associated with the misuse of MGC. [3] conducted a survey on the applications of MGC, reporting that it may contribute to academic misconduct, the spread of fake political news, and an increase in social media spam. Our goal in this paper is the effective detection of document-level paraphrased MGC, which is rewritten by LLMs from an original human-written text.
Automatically distinguishing MGC from Human-Produced Content (HPC) is, however, very challenging, given that MGC tends to resemble HPC, often containing at least chunks that are reproduced from human-produced training data. Currently, methods to detect MGC involve a combination of processes, namely a Natural Language Understanding (NLU) module that extracts linguistic and contextual features, followed by a classification module which fits this to binary labels using algorithms such as Logistic Regression (LR) [4] for decision-making.
However, these suffer from various limitations. First, current NLU models built using Transformers [5] such as Roberta [6] demonstrate good performance in analysing short texts (such as sentences) but less so when processing document-level data, as shown in [7]. They tend to struggle with retaining information if the texts are long. Even LLMs are limited in the length of the input texts that they can process effectively. Document-level texts exhibit more complex structures’ relationships. Secondly, the degree of similarity between MGC and HPC is higher, especially when LLMs are used to paraphrase HPC rather than generating it from scratch. This makes distinguishing HPC and MGC challenging, as differences in linguistic content are less distinctive. Previous work focuses on identifying similarities at surface-level linguistic features, ignoring the importance of other implicit features, such as structural features, which we show can contribute to the effectiveness of a detection model. There is indeed an implicit difference in Figure 1: the left side is more poetically structured, while the right side is a plainer narrative, though the distinction may not be obvious to the human reader. These limitations lead to underperforming detectors in work to date.
We propose a novel approach to automatically detect MGC which addresses the limitations described above. The aim is to better classify the texts by identifying structural features such as discourse relations. Specifically, we introduce a new model: DTransformer, which incorporates discourse features to improve detection accuracy.
@@
  • We create paraphrased Long-Form Question and Answer (paraLFQA) and paraphrased Writing Prompts (paraWP) – two parallel HPC-MGC corpora for model evaluation and further research in Section III
  • We introduce DTransformer – a discourse-aware Transformer-based model to incorporate discourse features in detection, which leads to better performance for long-form texts in Section IV.
II Related work
II-A MGC Detection Approaches
@@
In the academic literature, two types of models are commonly used: statistical methods, also known as feature-based methods, and neural network methods. [8] identified seven linguistic features and criteria for distinguishing MGC from HPC, and they concluded that MGC tends to exhibit fewer discourse and epistemic words but displays greater lexical diversity and nominalizations. [9] and Generative Language Test Room (GLTR) [10] also made decisions with linguistic features. Moreover, syntactic, sentimental, grammatical, and readability features are used to train models for fake news detection [9]. Apart from those linguistic features, there is N-gram-based methods, along with LLMs, have also been proposed for detection without requiring additional training, as demonstrated by DNA-GPT [11].
@@
However, feature-based methods need manual work to design the features. There have been studies comparing neural network methods and statistical methods [12], demonstrating the superior efficiency of Transformer-based models [5], such as Roberta Large OpenAI Detector [13], the first fake news detector Grover [14], a zero-shot detection method DetectGpt [15], and watermarking-based detection [16]. The Roberta Large Detector has been fine-tuned with 1.5B outputs generated with GPT2. The detection is treated as a binary classification. However, it has not taken into account additional linguistic characteristics of the texts that could enhance overall performance. More importantly, it was trained on texts produced with GPT-2, which may retain biases from GPT-2, while current machine-generation methods have evolved (e. g., GPT-4).
@@
Additionally, Fu et al. [16] have introduced semantic-aware watermarking algorithms, while [17] employ perturbation strategies to evade detection methods based on watermarking. Besides, Hu et al. [18] introduce another adversarial learning approach named Radar, noted for its robust transferability. Moreover, [19] directly utilises LLMs [20] for MGC detection, but concludes that these models are currently inadequate. Conversely, [21] combine two LLMs, BERT [22], and T5 [23], to improve detection efficacy. Moreover, recent advancements include the M4 model for multi-modality MGC detection [24] and DeBERTa [25] for monitoring mainstream news websites [26].
@@
Some efforts have been made specially on paraphrased MGC detection. [27] and [28] proposed using pre-trained word embedding models for text representation, combined with classifiers, to predict paraphrased plagiarism in writing. Animesh et al. [29] introduced an adversarial method to detect paraphrasing by leveraging the inferential properties of sentences. Other research has focused on developing corpus-based methods for detecting paraphrasing [30]. However, paraphrasing involves more than just substituting words and should incorporate additional detectable features. Researchers have manually crafted features, as discussed by Shahmohammadi et al. [31], which align with syntactic similarity, and other factors. Unlike these methods, Li et al. [32] have developed a paraphrase detection approach for text spans. While simple, this method offers an alternative perspective.
Moreover, numerous widely utilized commercial tools include GPTZero [^2] https://gptzero.me/, Copyleaks [^3] https://copyleaks.com/ai-content-detector, Content at Scale [^4] https://contentatscale.ai/ai-content-detector/, Originality [^5] https://originality.ai/, Winston AI [^6] https://gowinston.ai/, and Hive [^7] https://hivemoderation.com/ai-generated-content-detection. However, these tools utilize models that are not publicly accessible, which makes understanding these tools challenging.
As discussed above, the open challenges that have not been covered in previous work are as follows: Combining implicit features with linguistic features, and detecting document-level paraphrased machine-generated texts.
II-B Detection of Document-level MGC
@@
Long texts have more complex language organization compared to shorter texts, so comprehending long texts is more challenging than shorter ones. Current transformer-based models can fail when processing longer sequences [33] and there remains a risk of information loss due to the constrained memory and computational costs associated with self-attention operations [34]. One example is that Krishna et al. [35] developed an attack method targeting recent detectors by employing paraphrased sentences in MGC, resulting in a significant decrease in detector accuracy. This approach utilized paraphrased sentence patterns introduced by an external model known as DIPPER [35], highlighting the challenges associated with detecting discourse-paraphrased sentences which combines LLMs such as T5 and a sentence shuffle strategy to generate training data for discourse paraphrasing with high semantic similarity. As a result, Krishna et al. discovered that discourse-paraphrased sentences are difficult to detect.
@@
To address the aforementioned issue, hierarchical models have been proposed that handle sentence-level information before passing on these representations for document-level processing. [36] proposed the HIBERT model for document summarization. Moreover, [34] divided long sequences into different blocks, while SPECTER [37] was proposed to embed research papers by utilizing only the metadata. Longformer [38] introduced a sparse attention mechanism that adapts to sequence length, and [39] proposed to globalize sequences processed by BERT using a bidirectional GRU network to propagate information. In-context learning in OUTFOX [40] has been adapted for training on long essays. However, it does not explicitly learn structural organization but is instead stimulated by certain contextual relationships. Each of these models aims to address memory limitations and computational complexity by selecting representative embeddings. However, there are no dominating models utilizing discourse features.
@@
The discourse features could be represented by two annotations: Rhetorical Structure Theory Discourse Treebank (RST-DT) [41], a dataset that captures discourse relations within text spans, represented in a tree structure and the Penn Discourse Treebank (PDTB) [42], offering a straightforward representation of relations, specifying two arguments along with their relations [43] (as depicted in Figure 2). Annotated by human experts, PDTB’s 3.0 version includes over 100 types of relations, demonstrating exceptional applications when combined with Bert models [44]. These annotated labels are detailed about the role of each sentence as well.
@@
Figure 2: PDTB annotation example [45]: It shows the implicit relations, ACCORDINGLY, between the given two sentence.
III Datasets
@@
Currently, limited public datasets are available that focus on document-level paraphrased texts for MGC detection. For example, the recently published adversarial perturbations dataset introduced revised MGC texts with word replacements [46], but the dataset does not include meaningful differences in writing style between the two versions. Its primary aim is to attack current detectors. It is difficult to adopt this dataset for our research. Two other publicly available datasets are the Plagbench [47] and Multi-domain, and Multi-lingual Black-Box Machine-Generated Text Detection (M4) dataset [48]. Further, in this work, we introduce two new dataset that we created based on a portion of well-known datasets used in other NLP tasks: paraLFQA based on LFQA [49], which was also used in the DIPPER dataset [35] and paraWP based on WritingPrompts (WP) [50].
@@
The Plagbench dataset [47] is generated by GPT specifically for the task of plagiarism detection. We select the paraphrasing plagiarism subset, which includes human evaluation based on specially designed criteria such as semantic equivalence, adding credibility to its quality. The dataset consists of 135 data pairs.
@@
The M4 dataset contains a diverse array of human-written sources, including document-level texts from WikiHow, and others. It also includes MGC versions produced by a wide range of generators, such as BLOOMz [51] and ChatGPT. These machine-generated texts often differ in meaning from the original sources because the prompts used are not direct paraphrases but involve expansions, such as the prompt “Write a peer review based on the article.” It has been extensively utilized as a benchmark in MGC detection research. The dataset has a total of 133,551 text samples.
@@
For the paraLFQA dataset, the original answers from LFQA are considered as HPC. MGC was produced using the DIPPER model from the HPCs, resulting in discourse-paraphrased-style sentences, which we present as a new dataset. In accordance with the methodology outlined in the original paper, we apply T5 as the tokenizer. We selected a lexicon code of 40 and a re-order code of 100. These values represent that 40% of the lexicon and 100% of the order have been modified. These hyperparameters were chosen to emphasize changes in the organization of the language within the dataset. The dataset has around 2,000 MGC-HPC paragraph pairs with around 250-350 words in each paragraph. This serves as a control dataset where the semantic features are similar between the paraphrased and original content, but the structure organisation is different. One example is shown in Table I. It demonstrates that the sentence structures are entirely different, while the overall meaning remains the same.
Original Case (HPC) Paraphrased Case (HPC)
Pushing the bulbous part does NOT create a vacuum, no. It just pushed the air in the balloon towards the end of the pipette. If you release your fingers and the balloon wants to return to it’s initial round form (we come to that later on again), then the fluid at the entrance gets pushed in. The atmosphere that is in contact with the fluid pushes the fluid into the pipette. It does this because the balloon has some tension in its surface, and it wants to be round and flat, and not squeezed together. So as it opens up, the additional volume created within the balloon is filled due to the pressure from the outside atmosphere pushing onto the surface of the fluid, which in turn gets pushed into the pipette Awesome question, BTW! First of all, you’re not creating a vacuum by pushing on the bulb. What you’re doing is pushing the air in the bulb towards the tip of the tube. Now, when you let go of the bulb and it wants to go back to its original shape (we’ll come back to that), the liquid in the tip will be pushed in. Basically, the liquid in the bulb has a certain tension to it and, being a bulb, wants to be flattened, not squeezed, so, when it’s flattened, it fills with air from outside and the liquid is pushed into the tube. A very good question.
TABLE I: One case for comparison between original (HPC) and discourse paraphrased (MGC) text in paraLFQA.
@@
For the paraWP dataset, unlike paraLFQA focusing especially on discourse paraphrase, paraWP perform a generally much longer document-level dataset than paraLFQA. Specifically, human-written stories from the provided prompts are considered as HPC. The MGC paraphrases are produced by GPT-3.5-turbo with the following prompt: “Summarize the text. Expand the idea into new text. Keep the original narrative, such as he, she, I, or their names. Return the same number of words as the input texts.” In this way, the dataset has one-to-one matched text pairs with HPC and MGC, which are regarded as paraphrased-style versions. We have created 166,247 documents both for HPC and MGC in the training set, and 12,800 and 12,316 documents for the validation and test sets, respectively. We used the BPE tokenizer to process the data. The vocabulary size of the BPE modelling is 8,016. We have also investigated linguistic features of the dataset in Appendix A.
We also selected a portion of the paraWP dataset for revision to introduce surface-level word twist attacks by simply replacing certain words. This has been shown to be challenging for other detectors [52]. We manually performed the revision task to generate a small number of modified paraWP cases by: (a) incorporating present or past participles as adverbial clauses and using appositives, and (b) employing metaphors and parallel structures in sentences. This approach systematically reduces the frequency of adverbial clauses with -ing and -ed forms, appositives, and parallel sentence structures, which are surface-level word twist attacks.
@@
IV Model: DTransformer
This section introduces our method, specifically designed to address the second challenge: detecting document-level MGC texts paraphrased from HPC. The DTransformer model focuses on integrating structural features with semantic features, aiming to detect MGC by identifying the ways humans and machines organise text blocks, such as paragraphs. This model is inspired by the discourse information extraction method with PDTB annotation and the effects of structural features discussed in the Multi-Method Qualitative Text Framework [53]. The model is depicted in Figure 3, with detailed sub-modules discussed in the following paragraphs.
Figure 3: DTransformer model: it incorporates both structural and semantic features. It first splits the paragraphs and employs a hierarchical model to capture document semantic features. Then, using a cross-attention mechanism, it analyses discourse features for improved classification.